# Event2Vec: Processing Neuromorphic Events directly by Representations in Vector Space

### v3

@@ Abstract
Neuromorphic event cameras possess superior temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, their asynchronous and sparse data format poses a significant challenge for conventional deep learning methods. Existing solutions to this incompatibility often sacrifice temporal resolution, require extensive pre-processing, and do not fully leverage GPU acceleration. Inspired by word-to-vector models, we draw an analogy between words and events to introduce event2vec, a novel representation that allows neural networks to process events directly. This approach is fully compatible with the parallel processing and self-supervised learning capabilities of Transformer architectures. We demonstrate the effectiveness of event2vec on the DVS Gesture, ASL-DVS, and DVS-Lip benchmarks. A comprehensive ablation study further analyzes our method’s features and contrasts them with existing representations. The experimental results show that event2vec is remarkably parameter-efficient, has high throughput, and can achieve high accuracy even with an extremely low number of events. Beyond its performance, the most significant contribution of event2vec is a new paradigm that enables neural networks to process event streams as if they were natural language. This paradigm shift paves the way for the native integration of event cameras with large language models and multimodal models. Code, model, and training logs are provided in https://github.com/Intelligent-Computing-Lab-Panda/event2vec.
@@ 1 Introduction
Most contemporary deep learning models are designed to operate on dense, regularly structured, multi-dimensional tensors. This tensor-based paradigm is foundational to mainstream deep learning (LeCun et al., 2015) and is ubiquitously employed in modern scientific computing and machine learning frameworks, including NumPy (Harris et al., 2020a), TensorFlow (Abadi et al., 2016), and PyTorch (Paszke et al., 2019). Consequently, the sparse and asynchronous nature of event streams in the AER format is fundamentally incompatible with these tensor-based methods. To address this disparity, substantial research efforts have been devoted to converting events to dense representations, or designing new data and network structures to process the irregular events directly.
Figure 1: Conceptual analogy between words and events. The illustration of the DVS 128 camera is adapted from Lichtsteiner et al. (2008).
Existing methods primarily address the challenge of event encoding: how to effectively extract information from events and represent it for processing by neural networks. This challenge is analogous to word encoding in natural language processing, a problem successfully addressed by foundational techniques such as word-to-vector (word2vec) (Mikolov et al., 2013). The word2vec model embeds each word into a fixed-length vector, enabling the relationships between words to be represented by mathematical operations between vectors. This vector representation approach is highly compatible with deep learning architectures and has become a foundational component of modern Natural Language Processing (NLP) models (Devlin et al., 2019a; Brown et al., 2020). We identify numerous parallels between words and events, as illustrated in Figure 1. The key similarities are as follows:
@@
    The meaning of an element is determined by its context. A word can be polysemous; for instance, “transformer” can refer to a neural network architecture or a character in an animated series; its specific meaning is disambiguated by the surrounding text. An individual event merely signals a brightness change at a specific pixel and time, conveying little information in isolation. However, when viewed within a spatiotemporal stream, a sequence of events can delineate an object’s contour, thus giving a single event a higher-level meaning, such as being part of an edge. Therefore, the significance of an event is also fundamentally context-dependent.
@@
    We propose a parametric spatial embedding and a convolution-based temporal embedding method that captures neighborhood similarity—a task that is critical for accuracy but difficult for a standard embedding layer to learn.
@@
    We validated our method on three classification benchmarks: DVS Gesture, ASL-DVS, and DVS-Lip. It achieved accuracy competitive with traditional methods while demonstrating remarkable parameter efficiency, throughput, and robustness, especially with a low number of events.
Beyond the performance metrics, the most significant contribution of event2vec is its ability to enable neural networks to process event streams in a manner analogous to natural language. Therefore, state-of-the-art NLP architectures and methods—such as Transformer variants, model acceleration algorithms, and generative self-supervised training—can be directly leveraged for event-based vision. Event2vec paves the way for the native integration of event cameras with large language models and multimodal models.
@@ 2 Related Work
2.1 Dense Representations of events
Dense representations, derived from raw event streams, are fully compatible with conventional deep learning methods. This is typically achieved by integrating events along the time axis to form dense 3D or 4D tensors, such as event frames (Liu & Delbruck, 2018), multi-channel images (Barchid et al., 2022), voxel grids (Bardow et al., 2016), volumetric cubes (Cordone et al., 2022) and patches (Sabater et al., 2023; Peng et al., 2023).
@@ 2.1 Dense Representations of events
Specifically, event-to-frame methods accumulate events within discrete time intervals. These resulting frames can then be processed directly by standard neural networks. However, a significant drawback of these methods is the degradation or complete loss of the high temporal resolution inherent to event data. This occurs because individual event timestamps are aggregated or quantized during the conversion process. Furthermore, transforming the data into a dense representation negates the inherent spatial sparsity of events. For instance, the generated frames often contain a substantial number of zero-valued pixels. These pixels, while carrying no information, still incur significant memory and computational overhead. While many methods use timestamps implicitly to define the integration interval, some approaches explicitly leverage them to generate temporal weights (Zhu et al., 2019; Gehrig et al., 2019). Finally, the conversion process itself can be computationally intensive, introducing considerable latency that is often prohibitive for real-time applications (Rebecq et al., 2019; Gallego et al., 2022).
2.2 Irregular Representations of events
@@
Conversely, methods for processing irregular representations aim to preserve the inherent sparsity and asynchronicity of event data. This category includes SNNs (Maass, 1997; Roy et al., 2019), Sparse Convolutional Networks (Sparse CNNs) (Messikommer et al., 2020; Santambrogio et al., 2024), Graph Neural Networks (GNNs) (Bi et al., 2019; Schaefer et al., 2022), and point-based methods (Yang et al., 2019; Sekikawa et al., 2019; Lin et al., 2023; Ren et al., 2025).
@@
When deployed on neuromorphic hardware (Merolla et al., 2014; Davies et al., 2018), Spiking Neural Networks (SNNs) can process events in a naturally asynchronous event-driven manner. However, on standard hardware, GPU-based simulations of SNNs produce dense tensor outputs, as the hardware necessitates synchronous processing with discrete time-steps. Consequently, training SNNs on conventional GPUs typically occurs in a synchronous fashion, leading to an unavoidable performance gap between synchronous training and asynchronous inference (Yao et al., 2024; Du et al., 2025). Moreover, the reliance on backpropagation-through-time renders the training process slow and memory-intensive. Sparse CNNs leverage the inherent sparsity of event data, achieving a theoretically low number of Floating-Point Operations (FLOPs). Nevertheless, the architecture of standard GPUs is not optimized for the dynamic computations and unstructured memory access patterns required for efficient sparse acceleration. Consequently, similar to SNNs, Sparse CNNs fail to fully exploit the massive parallel processing capabilities of GPUs.
@@
Event-based GNNs construct graphs from incoming events, an approach that effectively preserves the spatio-temporal relationships between them. Since empty regions with no event activity do not generate graph nodes, the data’s sparsity is well-utilized. Their main disadvantage lies in the need for careful hyper-parameter tuning, such as the event downsampling rate and neighborhood radius for graph construction. Additionally, functioning as low-pass filters (Nt & Maehara, 2019), GNNs are susceptible to the over-smoothing problem (Zhou et al., 2020), which limits their ability to form deep architectures comparable to modern CNNs and Transformers (Vaswani et al., 2017). Point-based methods treat events from event cameras as analogous to point clouds from Light Detection and Ranging (LiDAR) sensors. A fundamental limitation of most point cloud models is their permutation invariance, which necessitates treating the input as an unordered set. Consequently, the event timestamp is typically relegated to being an additional positional coordinate, thereby discarding the crucial causal ordering of events. To manage the data volume, these methods often employ classic point cloud pre-processing techniques like farthest point sampling, which further increases latency.
@@ 3.1 Representing Events in a Vector Space
where $\textbf{v}\in\mathbb{R}^{D}$ is the resulting $D$-dimensional embedding vector, $\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)\in\mathbb{R}^{D}$ is the spatial embedding module, and $\textbf{v}_{t}=\text{Embed}_{t}(t)\in\mathbb{R}^{D}$ is the temporal embedding module. As shown in Eq. 1, this method fuses spatial and temporal information through addition. This additive fusion strategy is directly inspired by the positional encoding mechanism prevalent in Transformers.
@@ 3.2 Spatial Embedding
$\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)=\textbf{W}_{s}[p\cdot H\cdot W+y\cdot W+x],$ (2)
@@
However, this standard embedding layer imposes no inductive bias on the relationship between indices, compelling the model to learn all spatial relationships from data alone. In a tokenizer, a word’s index is a non-semantic identifier, whose order is primarily determined by the token’s frequency in the training corpus. Consequently, the words at indices $i$ and $i+1$ share no inherent semantic similarity. This assumption does not hold for event coordinates. Images are continuous two-dimensional functions (Gonzalez, 2009). Spatially adjacent pixels are known to exhibit strong correlation. Therefore, an effective spatial embedding should incorporate this locality bias, ensuring that events with close coordinates yield similar embedding vectors:
$\text{Embed}_{s}(x+\Delta x,y+\Delta y,p)-\text{Embed}_{s}(x,y,p)\approx\textbf{0},$ (3)
for small coordinate perturbations $\Delta x,\Delta y]$, e.g., $\Delta x,\Delta y]=[1,0]$.
The standard embedding in Eq. 2 fails to account for this crucial spatial relationship, which can impede the learning process. To solve this issue, we propose an elegant parametric algorithm for generating the embedding matrix $\textbf{W}_{\phi}$ by a neural network $\phi$. To systematically enumerate all spatial coordinates within a $P\times H\times W$ volume (where $P=2$ represents the two polarities), we first establish a linear index sequence $\textbf{c}=[0,1,\dots,P\cdot H\cdot W-1]$. This sequence is then decomposed into three probe tensors, $\textbf{x}_{c}$, $\textbf{y}_{c}$, and $\textbf{p}_{c}$, which correspond to the coordinates along the width, height, and polarity dimensions, respectively. The transformation is defined as follows: $\textbf{x}_{c}=\textbf{c}\pmod{W},\textbf{y}_{c}=\left\lfloor\frac{\textbf{c}}{W}\right\rfloor\pmod{H},\textbf{p}_{c}=\left\lfloor\frac{\textbf{c}}{WH}\right\rfloor$. Finally, these probe tensors are passed through $\phi$, which outputs the complete embedding matrix $\textbf{W}_{\phi}=\phi(\textbf{x}_{c},\textbf{y}_{c},\textbf{p}_{c})$. By substituting the parametrically generated matrix $\textbf{W}_{\phi}$ into the look-up mechanism of Eq. 2, we establish a direct equivalence for any given event coordinate $(x,y,p)$:
@@
$\phi(x+\Delta x,y+\Delta y,p)-\phi(x,y,p) \approx\frac{\partial\phi}{\partial x}\Delta x+\frac{\partial\phi}{\partial y}\Delta y+o(\|\Delta\|),$ (5)
where $o(\|\Delta\|)$ represents higher-order remainder terms. As Eq. 5 illustrates, for small perturbations $\Delta x,\Delta y]$, the difference between the embeddings is proportional to the gradient of $\phi$. Consequently, as the perturbations approach zero, this difference vector also approaches zero:
$\phi(x+\Delta x,y+\Delta y,p)-\phi(x,y,p)\approx\textbf{0}.$ (6)
In this manner, the use of a continuous parametric network $\phi$ inherently embeds the desired neighborhood semantics, or spatial inductive bias, directly into the embedding matrix. This approach elegantly satisfies the condition outlined in Eq. 3.
@@ 3.3 Temporal Embedding
However, directly applying these relative positional encoding techniques to event timestamps is ill-suited. Such methods are fundamentally designed for discrete and uniformly spaced indices, whereas event timestamps are continuous and inherently non-uniform. To address this discrepancy, we propose learning the temporal embedding directly from the differences between consecutive timestamps using a neural network.
Specifically, the temporal embedding module is implemented as a stack of convolutional layers. For the $i$-th event in an event stream, the input to this module is the first-order temporal difference, t[i+1]t[i]t[i+1]-t[i]. This design offers several distinct advantages:
@@
    Time-Shift Invariance: By operating on relative temporal distances, the embedding becomes inherently invariant to absolute shifts in time.
@@
    Generalization: It functions as a variant of position-wise learnable encoding but circumvents the length generalization problem by accepting continuous values as input, rather than being tied to a fixed vocabulary of discrete positions.
@@
    Contextual Consistency: The convolutional operations allow the temporal embedding for an event to be influenced by the timing of its immediate neighbors, thereby reinforcing the principle of neighborhood semantics in the time domain.

### v4

@@ Abstract
Neuromorphic event cameras possess superior temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, their asynchronous and sparse data format poses a significant challenge for conventional deep learning methods. Existing methods either convert the events into dense synchronous frame representations for processing by powerful CNNs or Transformers, but lose the asynchronous, sparse and high temporal resolution characteristics of events during the conversion process; or adopt irregular models such as sparse convolution, spiking neural networks, or graph neural networks to process the irregular event representations but fail to take full advantage of GPU acceleration. Inspired by word-to-vector models, we draw an analogy between words and events to introduce event2vec, a novel representation that allows neural networks to process events directly. This approach is fully compatible with the parallel processing capabilities of Transformers. We demonstrate the effectiveness of event2vec on the DVS Gesture, ASL-DVS, and DVS-Lip benchmarks, showing that event2vec is remarkably parameter-efficient, features high throughput and low latency, and achieves high accuracy even with an extremely low number of events or low spatial resolutions. Event2vec introduces a novel paradigm by demonstrating for the first time that sparse, irregular event data can be directly integrated into high-throughput Transformer architectures. This breakthrough resolves the long-standing conflict between maintaining data sparsity and maximizing GPU efficiency, offering a promising balance for real-time, low-latency neuromorphic vision tasks. The code is provided in https://github.com/Intelligent-Computing-Lab-Panda/event2vec.
Neuromorphic Computing, Event Camera
@@ 1 Introduction
Most contemporary deep learning models are designed to operate on dense, regularly structured, multi-dimensional tensors. This regular paradigm is foundational to mainstream deep learning (LeCun et al., 2015) and is ubiquitously employed in modern scientific computing and machine learning frameworks, including NumPy (Harris et al., 2020), TensorFlow (Abadi et al., 2016), and PyTorch (Paszke et al., 2019). Consequently, the sparse and asynchronous nature of event streams in the AER format is fundamentally incompatible with these regular methods. To address this disparity, substantial research efforts have been devoted to converting events to dense representations, or designing new data and network structures to process the irregular events directly.
Figure 1: Conceptual analogy between words and events.
Existing methods primarily address the challenge of event encoding: how to effectively extract information from events and represent it for processing by neural networks. This challenge is analogous to word encoding in natural language processing, a problem successfully addressed by word-to-vector (word2vec) (Mikolov et al., 2013). The word2vec model embeds each word into a fixed-length vector, enabling the relationships between words to be represented by mathematical operations between vectors. This vector representation approach is highly compatible with deep learning architectures and has become a foundational component of modern Natural Language Processing (NLP) models (Devlin et al., 2019a; Brown et al., 2020). As illustrated in Figure 1, we identify numerous parallels between words and events. The key similarities are as follows:
@@
    The meaning of an element is determined by its context. A word can be polysemous; for instance, “transformer” can refer to a neural network architecture or a character in an animated series; its specific meaning is disambiguated by the surrounding text. An individual event merely indicates a brightness change at a specific pixel and time, conveying little information in isolation. However, when viewed within a spatiotemporal stream, a sequence of events can delineate an object’s contour, thus giving a single event a higher-level meaning, such as being part of an edge. Therefore, the significance of an event is also fundamentally context-dependent.
@@
    We propose a parametric spatial embedding and a convolution-based temporal embedding method to capture neighborhood similarity—a characteristic that is critical for accuracy but difficult for a standard embedding layer (a look-up table) to learn.
@@
    We validated our method on three widely used classification benchmarks: DVS Gesture (Amir et al., 2017), ASL-DVS (Bi et al., 2019), and DVS-Lip (Tan et al., 2022). It achieved competitive accuracy while demonstrating remarkable parameter efficiency, throughput, latency, as well as robustness against a low number of events or low spatial resolutions.
Existing methods either leverage CNNs or Transformers to process dense representations of events, which exploit extremely high computational efficiency on GPUs and enable large-scale model deployment, but sacrifice the sparse, asynchronous, and high temporal resolution characteristics inherent to events; or adopt irregular network architectures (e.g., sparse convolution, spiking neural networks, graph neural networks) to handle raw or downsampled sparse representations of events, yet these irregular models can hardly take full advantage of the massive parallel computing capability of GPUs. Our proposed method, event2vec, addresses this dilemma for the first time, allowing the simultaneous use of sparse irregular representations and dense regular models, and thus provides a brand-new event processing paradigm for neuromorphic vision.
@@ 2 Related Work
2.1 Dense Representations and Processing of Events
@@
Dense representations, derived from raw event streams, are fully compatible with conventional deep learning methods. This is typically achieved by integrating events along the time axis to form dense 3D or 4D tensors, such as event frames (Liu and Delbruck, 2018), multi-channel images (Barchid et al., 2022), voxel grids (Bardow et al., 2016), volumetric cubes (Cordone et al., 2022), and patches (Sabater et al., 2023; Peng et al., 2023). Specifically, event-to-frame methods accumulate events within discrete time intervals. The resulting frames can then be processed directly by standard neural networks. However, a significant drawback of these methods is the degradation of the high temporal resolution inherent to event data. This occurs because individual event timestamps are aggregated or quantized during the conversion process. Furthermore, transforming the data into a dense representation negates the inherent spatial sparsity of events. For instance, the generated frames often contain a substantial number of zero-valued pixels. These pixels, while carrying no information, still incur significant memory and computational overhead. While many methods use timestamps implicitly to define the integration interval, some approaches explicitly leverage them to generate temporal weights (Zhu et al., 2019; Gehrig et al., 2019). Finally, the conversion process itself can be computationally intensive, introducing considerable latency that is often prohibitive for real-time applications (Rebecq et al., 2019; Gallego et al., 2022).
2.2 Irregular Representations and Processing of Events
@@
Conversely, methods for processing irregular representations aim to preserve the inherent sparsity and asynchronicity of event data. This category includes Spiking Neural Networks (SNNs) (Maass, 1997; Roy et al., 2019), Sparse Convolutional Networks (Sparse CNNs) (Messikommer et al., 2020; Santambrogio et al., 2024), Graph Neural Networks (GNNs) (Bi et al., 2019; Schaefer et al., 2022), and point-based methods (Yang et al., 2019; Sekikawa et al., 2019; Lin et al., 2023; Ren et al., 2025).
@@
When deployed on neuromorphic hardware (Merolla et al., 2014; Davies et al., 2018), SNNs can process events in a naturally asynchronous, event-driven manner. However, on standard hardware, GPU-based simulations of SNNs produce dense tensor outputs, as the hardware necessitates synchronous processing with discrete time-steps. Consequently, training SNNs on GPUs typically occurs in a synchronous fashion, leading to an unavoidable performance gap between synchronous training and asynchronous inference (Yao et al., 2024; Du et al., 2025). Moreover, the reliance on backpropagation-through-time renders the training process slow and memory-intensive. Sparse CNNs leverage the inherent sparsity of event data, achieving a theoretically low number of Floating-Point Operations (FLOPs). Nevertheless, the architecture of GPUs is not optimized for the dynamic computations and unstructured memory access patterns required for efficient sparse acceleration. Consequently, similar to SNNs, Sparse CNNs fail to fully exploit the massive parallel processing capabilities of GPUs.
@@
Event-based GNNs construct graphs from incoming events, an approach that effectively preserves the spatio-temporal relationships between them. Since empty regions with no event activity do not generate graph nodes, the data’s sparsity is well-utilized. Their main disadvantage lies in the need for careful hyper-parameter tuning, such as the event downsampling rate and neighborhood radius for graph construction. Additionally, functioning as low-pass filters (Nt and Maehara, 2019), GNNs are susceptible to the over-smoothing problem (Zhou et al., 2020), which limits their ability to form deep architectures comparable to modern CNNs and Transformers (Vaswani et al., 2017). Point-based methods treat events from event cameras as analogous to point clouds from Light Detection and Ranging (LiDAR) sensors. A fundamental limitation of most point cloud models is their permutation invariance, which necessitates treating the input as an unordered set. Consequently, the event timestamp is typically relegated to being an additional positional coordinate, thereby discarding the crucial causal ordering of events. To manage the data volume, these methods often employ classic point cloud pre-processing techniques like farthest point sampling, which further increases latency.
@@ 3.1 Representing Events in a Vector Space
where $\textbf{v}\in\mathbb{R}^{D}$ is the resulting $D$-dimensional embedding vector, $\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)\in\mathbb{R}^{D}$ is the spatial embedding vector, and $\textbf{v}_{t}=\text{Embed}_{t}(t)\in\mathbb{R}^{D}$ is the temporal embedding vector. As shown in Eq. 1, this method fuses spatial and temporal information through addition. This additive fusion strategy is directly inspired by the positional encoding mechanism prevalent in Transformers.
@@ 3.2 Spatial Embedding
$\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)=\textbf{W}_{s}[p\cdot H\cdot W+y\cdot W+x],$
@@
However, this standard embedding layer imposes no inductive bias on the relationship between indices, compelling the model to learn all spatial relationships from data alone. In a tokenizer, a word’s index is a non-semantic identifier, the assignment of which is primarily determined by the word’s frequency in the training corpus. Consequently, the words at indices $i$ and $i+1$ share no inherent semantic similarity. This assumption does not hold for event coordinates. Images are continuous two-dimensional functions (Gonzalez, 2009). Spatially adjacent pixels are known to exhibit strong correlation. Therefore, an effective spatial embedding should incorporate this locality bias, ensuring that events with close coordinates yield similar embedding vectors:
$\text{Embed}_{s}(x+\Delta x,y+\Delta y,p)-\text{Embed}_{s}(x,y,p)\approx\textbf{0},$
for small coordinate perturbations $(\Delta x,\Delta y)$.
The standard embedding in Eq. 2 fails to account for this crucial spatial relationship, which can impede the learning process. To solve this issue, we propose an elegant parametric algorithm for generating the embedding matrix $\textbf{W}_{\phi}$ via a neural network $\phi$. To systematically enumerate all spatial coordinates within a $P\times H\times W$ volume (where $P=2$ represents the two polarities), we first establish a linear index sequence $\textbf{c}=[0,1,\dots,P\cdot H\cdot W-1]$. This sequence is then decomposed into three probe tensors, $\textbf{x}_{c}$, $\textbf{y}_{c}$, and $\textbf{p}_{c}$, which correspond to the coordinates along the width, height, and polarity dimensions, respectively. The transformation is defined as follows: $\textbf{x}_{c}=\textbf{c}\pmod{W},\textbf{y}_{c}=\left\lfloor\frac{\textbf{c}}{W}\right\rfloor\pmod{H},\textbf{p}_{c}=\left\lfloor\frac{\textbf{c}}{WH}\right\rfloor$. Finally, these probe tensors are passed through $\phi$, which generates the complete embedding matrix $\textbf{W}_{\phi}=\phi(\textbf{x}_{c},\textbf{y}_{c},\textbf{p}_{c})$. By substituting the parametrically generated matrix $\textbf{W}_{\phi}$ into the look-up mechanism of Eq. 2, we establish a direct equivalence for any given event coordinate $(x,y,p)$:
@@
$\phi(x+\Delta x,y+\Delta y,p)-\phi(x,y,p)$
$\approx\frac{\partial\phi}{\partial x}\Delta x+\frac{\partial\phi}{\partial y}\Delta y+o(\|\Delta\|),$ (5)
where $o(\|\Delta\|)$ represents higher-order remainder terms. As Eq. 5 illustrates, for small perturbations $(\Delta x,\Delta y)$, the difference between the embeddings is approximated by the inner product of the gradient of $\phi$ and the perturbation vector. Consequently, as the perturbations approach zero, this difference vector also approaches zero. In this manner, a continuous parametric network $\phi$ inherently embeds the desired neighborhood semantics, or spatial inductive bias, directly into the embedding matrix. This approach elegantly satisfies the condition outlined in Eq. 3.
@@ 3.3 Temporal Embedding
However, directly applying these relative positional encoding techniques to event timestamps is ill-suited. Such methods are fundamentally designed for discrete and uniformly spaced indices, whereas event timestamps are continuous and inherently non-uniform. To address this discrepancy, we propose learning the temporal embedding directly from the differences between consecutive timestamps.
Specifically, the temporal embedding module is implemented as a stack of convolutional layers, which take the sequence of the first-order temporal difference of timestamps $\Delta\textbf{t}=[\textbf{t}[1]-\textbf{t}[0],\textbf{t}[2]-\textbf{t}[1],...,\textbf{t}[L-1]-\textbf{t}[L-2],0]$ as the input, where $L$ is the number of events, and the last 0 is the padding value. This design offers several advantages:
@@
    Time-Shift Invariance: By operating on relative temporal differences, the embedding becomes inherently invariant to absolute shifts in time.
@@
    Contextual Consistency: The convolutional operations allow the temporal embedding for an event to be influenced by the timing of its immediate neighbors, thereby reinforcing the principle of neighborhood semantics in the time domain. On the other hand, the occurrence of individual events may contain a certain amount of noise, and convolution is applied to achieve the effect of local smoothing and noise reduction.
@@
    Optimization Efficiency and Inductive Bias: Providing $\Delta\textbf{t}$ as input serves as a form of temporal ‘preconditioning’, aligning with the principle of residual learning (He et al., 2016). While a network could theoretically infer intervals from absolute timestamps t, explicitly modeling $\Delta\textbf{t}$ reduces the optimization burden by providing a direct representation of event velocity. Since the convolution operation essentially performs a weighted summation, it is mathematically congruent with differential inputs $\Delta\textbf{t}$. The summation of consecutive time differences possesses a clear physical interpretation: it represents the accumulated duration of a local event window, enabling the network to directly measure the local event density.